xx ExternalLibraries/yaml-cpp

ExternalLibraries/yaml-cpp

Roland Haas

Wed Jul 10 15:07:16 CEST 2019

Abstract

Thorn yaml-cppxprovides access to the yaml-cpp mesh refinement library.

1 Parameters

2 Interfaces General Implements: yamlxcpp

3 Schedule This section lists all the variables which are assigned storage by thorn ExternalLibraries/yamlxcpp. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE